# Robust certification of quantum instruments through a sequential communication game

### v1

@@ IV Sequential communication game in higher dimensions
Our proposed two-receiver communication game with restricted collaboration can be generalized to more sophisticated scenarios of QRAC with higher-dimensional systems and longer messages to be encoded or transmitted [21, 73]. Here we discuss the higher-dimensional generalization and compute numerically the lower bound of the optimal success probability employing the see-saw method up to dimension six (see Appendix C for the details). The see-saw is an iterative optimization method for affine functions of Hermitian operators, which was introduced initially to maximize quantum violation of a given Bell inequality [74].
The results are displayed in Fig. 4 where we separately show the success probalilities corresponding to Aparna, Barun and Chhanda, respectively, in the three sub-figures. In each of the sub-figures, we plot the respective success probabilities for three cases: (i) the classical success probability $P^{C}$, (ii) the quantum success probability at the threshold of quantum advantage $P(\eta=\eta_{c})$, with $\eta_{c}$ denoting the critical value at which Barun’s success probability matches the classical benchmark, given by $\eta_{c}=\frac{d-1}{d+\sqrt{d}-2}$, and (iii) the quantum success probability for sharp measurement by Barun, $P(\eta=1)$. The magnitudes of all the success probabilities are seen to drop with an increase in dimensions. Note from Figs. 4(a) and 4(b) that while Barun’s success probability drops with the increase of the sharpness parameter, the reverse occurs in the case of Chhanda’s success probability. Thus, the trade-off between Barun’s and Chhanda’s success is clearly maintained even in higher dimensions. From Fig. 4(c) we observe that the magnitude of quantum advantage increases with the dimension. Moreover, the resourcefulness of the unsharp measurement in the context of the present communication task is obvious, outperforming its sharp counterpart.
@@ V Conclusions
Certification of quantum instruments [75] is a crucial ingredient of communication protocols relevant to quantum technology. Since noise is ubiquitous in most practical applications, the enhancement of robustness is an essential goal in such tasks. In the present work, we have proposed a new communication game extending the standard random access code paradigm to the sequential framework [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 59, 60, 61], involving one sender and two receivers with restricted communication allowed between the latter parties. Our protocol is based on sequential measurements in the prepare-transform-measure scenario, which has received much attention in recent times [31, 32, 33, 34, 65, 66, 67]. Through our proposed communication game, we can certify the sender’s state preparation and the measurement devices of the two sequential receivers.
@@
It may be worth reemphasizing that quantum random access codes are fundamental communication tasks, which can exhibit quantum advantage even for the simplest quantum system—a qubit. Such advantage is one of the central issues in quantum information science, facilitating many quantum information processing tasks [23, 24, 25, 26, 27, 28, 29, 30], and is successful in addressing certain fundamental questions as well [37, 38, 39, 40, 41]. Non-trivial extensions of the standard QRAC enable self-testing of quantum instruments and non-projective measurements in the presence of noise in realistic experimental scenarios [43, 44, 45, 68]. Here we have proposed a novel three-party communication game based on $2\rightarrow 1$ QRAC, which involves collaboration among the receivers, and is hence, fundamentally different from the earlier extensions.
@@
To summarize, in our proposed game, Aparna, the sender, encodes a two-bit message into a bit, and the task of the first receiver, Barun, is to decode a bit value according to his random input, whereas the second receiver, Chhanda, has to work out the leftover message. Thus, Barun and Chhanda may have some non-zero probability of correctly retrieving the full message, and they are allowed to have limited communication: Barun passes his input information to Chhanda, but there is no leakage of output information. Analyzing the protocol, we find a quantum advantage, which can certify Aparna’s state preparation, Barun’s instruments, as well as Chhanda’s measurement. Remarkably, the semi-device-independent certification of the sharpness parameter associated with Barun’s instrument is more robust in our protocol compared to that in the method based on sequential QRAC without any collaboration among the receivers [65]. We further generalize our sequential protocol to higher dimensions. By computing numerically the lower bound on the quantum success probability, we have shown that the quantum advantage increases with system dimensions, highlighting again the role of unsharp measurement in communication tasks relevant for sequential networks.
@@
The analysis presented here gives rise to prospects for future work in several possible directions. It would be interesting to see how robust the certification task based on the proposed quantum access code under restricted communication is in higher dimensions. Further, it would be worthwhile investigating implications of our proposed game in other RAC-based tasks of foundational and information processing relevance [23, 24, 25, 26, 27, 28, 30, 37, 38, 39, 40, 41]. Moreover, our present analysis should motivate explorations of more sophisticated generalizations of the RAC paradigm in the sequential measurement framework, such as in multipartite scenarios involving more than two receivers, towards the goal of harnessing even greater quantum advantage with minimal resources.
@@ Appendix A Optimal of success probability and self-testing
$M_{b|y}=f_{yb}\ket{\vec{g}_{y}}\bra{\vec{g}_{y}}+h_{yb}\ket{-\vec{g}_{y}}\bra{-\vec{g}_{y}}$ | (19)
Here, |$\ket{\vec{g}_{y}}$ is the pure state corresponding to the bloch sphere direction $\vec{g}_{y}$,
@@
$T=\sum_{y,b}\frac{|\vec{q}_{\bar{y}}|}{2}\sqrt{(1+(-1)^{b}g_{y0})^{2}-|\vec{g}_{y}|^{2}(1-\bra{\vec{g}_{y}}\hat{q}_{\bar{y}}\cdot\vec{\sigma}\ket{\vec{g}_{y}}^{2})}$ (22)
We now perform the optimization over $g_{y0}$ by considering the two terms corresponding to $y=0,1$ separately. This amounts maximizing an expression of the form $\sqrt{(1+w)^{2}-z}+\sqrt{(1-w)^{2}-z}$, for positive $z$. It can be shown such that such a function is uniquely maximized by setting $w=0$. So we require $g_{00}=g_{10}0$. Moreover, $(\vec{q}_{0},\vec{q}_{1})$ has no component along y-axis, and one can optimally choose $g_{02}=g_{12}=0$. This enables us to express
@@
Lemma 1: For every tuple $(\theta,\zeta_{0},\zeta_{1})$ corresponding to $(\beta,\tilde{P})$, there exists another tuple $(\theta,\zeta_{0},\zeta_{1})=(\frac{\pi}{2},\zeta,\zeta)$ that produces $(\beta,\tilde{P}^{{}^{\prime}})$ with $\tilde{P}^{{}^{\prime}}\geq\tilde{P}$, Moreover, $\theta=\frac{\pi}{2}$ and $\zeta_{0}=\zeta_{1}$ leads to the optimal $\tilde{P}^{{}^{\prime}}$.
@@
In order to analyze the self-testing argument explicitly, let us more closely examine the steps needed to arrive at Eq.  (30). First, we have already shown that the preparations must be pure, pairwise antipodal and by lemma 1, they must have a relative angle of $\frac{\pi}{2}$. This corresponds to a square in a disk of the Bloch sphere. The above arguments fully characterize Aparna’s preparations up to a reference frame. For Barun’s instrument, we have shown that the Bloch vectors ($\vec{g}_{0},\vec{g}_{1}$) can only have non-zero components in the $x$- and $z$-directions respectively, and that the length of the Bloch vector is given by Eq. (29). This fully characterizes the Bloch vectors. Moreover, in Eq. (15), we require that $D_{yb}^{\bar{y}}$ is aligned with the eigenvector of $\sqrt{M_{b|y}}D_{yb}\sqrt{M_{b|y}}$ corresponding to the largest eigenvalue. In case of Chhanda’s measurement, we now have $\gamma_{0}=\sigma_{x}$ and $\gamma_{1}=\sigma_{z}$, while $M_{0}\propto\sigma_{x}$ and $M_{1}\propto\sigma_{z}$. Therefore, it follows that $\forall y,b:D_{yb}^{0}=|+\rangle\langle+|$ and $\forall y,b:D_{yb}^{1}=|0\rangle\langle 0|$, we have,
$U_{yb}^{\dagger}N_{0|0}U_{yb}=\ket{+}\bra{+}$ | (31)
$U_{yb}^{\dagger}N_{0|1}U_{yb}=\ket{0}\bra{0}$ |
@@
Appendix B Upper and lower bound on the sharpness parameter $\eta$
@@
$\eta\geq\sqrt{2}(2P_{AB}^{Q}-1)$ (34)
For obtaining the upper bound of $\eta$, we make use of the optimal choice of the success probability from Eq.9, given
$P_{AC}^{Q}\leq\frac{1}{2}+\frac{\sqrt{1-\eta^{2}}}{2\sqrt{2}}$ (35)
Inverting the above inequality, leads to,
$\eta\leq\sqrt{8P_{AC}^{Q}-8(P_{AC}^{Q})^{2}-1}$ (36)
Appendix C Success probabilities for higher dimensional systems

### v2

@@ IV Sequential communication game in higher dimensions
Our proposed two-receiver new communication game can be generalized to more sophisticated scenarios of QRAC with higher-dimensional systems and longer messages to be encoded or transmitted [23, 75]. Here we discuss the higher-dimensional generalization and compute numerically the lower bound of the optimal success probability employing the see-saw method up to dimension six (see Appendix C for the details). The see-saw is an iterative optimization method for affine functions of Hermitian operators, which was introduced initially to maximize payoff function of the quantum strategy of a given Bell-type game [76].
The results are displayed in Fig. 4 where we separately show the individual payoff functions corresponding Barun and Chhanda and their joint payoff function as described in Section II in the three sub-figures. In each of the sub-figures, we plot the respective success probabilities for three cases: (i) the classical success probability $P^{C}$, (ii) the quantum success probability at the threshold of quantum advantage $P(\eta=\eta_{c})$, with $\eta_{c}$ denoting the critical value at which Barun’s success probability matches the classical benchmark, given by $\eta_{c}=\frac{d-1}{d+\sqrt{d}-2}$, and (iii) the quantum success probability for sharp measurement by Barun, $P(\eta=1)$. The magnitudes of all the success probabilities are shown to drop with an increase in dimensions whereas quantum advantage is increasing. Note from Figs. 4(a) and 4(b) that while Barun’s success probability drops with the increase of the sharpness parameter, the reverse occurs in the case of Chhanda’s success probability. Thus, the trade-off between Barun’s and Chhanda’s success is clearly maintained even in higher dimensions. From Fig. 4(c) we observe that the magnitude of quantum advantage increases with the dimension. Moreover, the resourcefulness of the unsharp measurement in the context of the present communication task is obvious, outperforming its sharp counterpart.
@@ V Conclusions
Certification schemes for various quantum components is essential in the context of present-day quantum science and technology, as malicious providers of devices may be a threat to the security and proper functioning of quantum protocols. For certification of quantum states and measurements, there are several schemes in the literature. for example, device-independent self-testing [4, 5, 6, 7, 8, 9, 10], semi device-independent certification [11, 12, 13, 14, 15, 16] and device dependent protocols [17, 18]. It is to be noted that there exist quantum components other than states and measurements which are not certified by the standard above mentioned certification schemes. Recently, sequential extension of the standard protocols has been proposed which can be employed to self-test quantum instruments [67], non-projective [69] and unsharp measurements [68] in a semi-device-independent manner going beyond the standard settings of certification of quantum states and measurements.
@@
It may be worth reemphasizing that quantum random access codes are fundamental communication tasks, which can exhibit quantum advantage even for the simplest quantum system — for a qubit. Such advantage is one of the central issues in quantum information science, facilitating many quantum information processing tasks [25, 26, 27, 28, 29, 30, 31, 32], and is successful in addressing certain fundamental questions as well [39, 40, 41, 42, 43]. Non-trivial extensions of the standard QRAC enable self-testing of quantum instruments and non-projective measurements in the presence of noise in realistic experimental scenarios [45, 46, 47, 69].
In the present work, we have proposed a new communication game extending the standard random access code paradigm to the sequential framework [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 61, 62, 63], involving one sender and two receivers with restricted communication allowed between the receivers. The novel three-party communication game is based on $2\rightarrow 1$ QRAC, which is fundamentally different from the earlier extensions. Our protocol is based on sequential measurements in the prepare-transform-measure scenario, which has received much attention in recent times [33, 34, 35, 36, 67, 68, 70]. We have shown prominent quantum advantage of the game and used it for the purpose of certification of relevant quantum components in a semi device-independent way and generalizes the game to the higher dimensional system as well.
@@
To summarize, in our proposed game, Aparna, the sender, encodes a two-bit message into a bit, and the task of the first receiver, Barun, is to decode a bit value according to his random input, whereas the second receiver, Chhanda, has to work out the leftover message. Thus, Barun and Chhanda may have some non-zero probability of correctly retrieving the full message, and they are allowed to have limited communication: Barun passes his post-measured state without leaking any output information. We have found a quantum advantage, which can certify Aparna’s state preparation, Barun’s instruments, as well as Chhanda’s measurement. Remarkably, the SDI certification of the sharpness parameter associated with Barun’s instrument is more robust in our protocol compared to that in the method based on sequential QRAC without any collaboration among the receivers [67]. We further generalize our sequential protocol to higher dimensions. By computing numerically employing seesaw algorithm, the lower bound on the quantum success probability, we have shown that the quantum advantage increases with system dimensions, highlighting again the role of unsharp measurement in communication tasks relevant for sequential networks.
@@
The analysis presented here gives rise to prospects for future work in several possible directions. It would be interesting to see how robust the certification task based on the proposed quantum access code under restricted communication is in higher dimensions. Further, it would be worthwhile investigating implications of our proposed game in other RAC-based tasks of foundational and information processing relevance [25, 26, 27, 28, 29, 30, 32, 39, 40, 41, 42, 43]. Moreover, our present analysis should motivate explorations of more sophisticated generalizations of the RAC paradigm in the sequential measurement framework, such as in multipartite scenarios involving more than two receivers, towards the goal of harnessing even greater quantum advantage with minimal resources. Finally, one may explore situations in which non-cooperation on the part of the first receiver is allowed by introducing cheating strategies modeled by adding noise or deforming the post-measurement state leading to reduction of the payoff functions.
@@ Appendix A Optimal of success probability and self-testing
[…]
@@
[…]
We now perform the optimization over $g_{y0}$ by considering the two terms corresponding to $y=0,1$ separately. This amounts maximizing an expression of the form $\sqrt{(1+w)^{2}-z}+\sqrt{(1-w)^{2}-z}$, for positive $z$. It can be shown such that such a function is uniquely maximized by setting $w=0$. So we require $g_{00}=g_{10}=0$. Moreover, $(\vec{q}_{0},\vec{q}_{1})$ has no component along y-axis, and one can optimally choose $g_{02}=g_{12}=0$. This enables us to express
@@
[…]
@@
In order to analyze the self-testing argument explicitly, let us more closely examine the steps needed to arrive at Eq.  (35). First, we have already shown that the preparations must be pure, pairwise antipodal, and by lemma 1, they must have a relative angle of $\frac{\pi}{2}$. This corresponds to a square in a disk of the Bloch sphere. The above arguments fully characterize Aparna’s preparations up to a reference frame. For Barun’s instrument, we have shown that the Bloch vectors ($\vec{g}_{0},\vec{g}_{1}$) can only have non-zero components in the $x$- and $z$-directions respectively, and that the length of the Bloch vector is given by Eq. (34). This fully characterizes the Bloch vectors. Moreover, in Eq. (20), we require that $D_{yb}^{\bar{y}}$ is aligned with the eigenvector of $\sqrt{M_{b|y}}D_{yb}\sqrt{M_{b|y}}$ corresponding to the largest eigenvalue. In case of Chhanda’s measurement, we now have $\gamma_{0}=\sigma_{x}$ and $\gamma_{1}=\sigma_{z}$, while $M_{0}\propto\sigma_{x}$ and $M_{1}\propto\sigma_{z}$. Therefore, it follows that $\forall y,b:D_{yb}^{0}=|+\rangle\langle+|$ and $\forall y,b:D_{yb}^{1}=|0\rangle\langle 0|$, we have,
$|\displaystyle U_{yb}^{\dagger}N_{0|0}U_{yb}=\ket{+}\bra{+}$ (36)
$|\displaystyle U_{yb}^{\dagger}N_{0|1}U_{yb}=\ket{0}\bra{0}$
@@
Appendix B Comparison of bound gaps for different visibility settings
@@
Appendix C Success probabilities for higher-dimensional systems